OpenStudy (anonymous):

If f()=-2x+3 and g(x)=4x-3, which is greater, f(5) or g(-2)?

OpenStudy (gorv):

f(5)=-2*5+3=-10+3=-7 g(-2)=4*(-2)-3=-8-3=-11

OpenStudy (gorv):

f(5) is greater

OpenStudy (anonymous):

thanks!